# An Empirical Study of Training ID-Agnostic Multi-modal Sequential Recommenders

### v2

@@
Abstract.
Sequential Recommendation (SR) aims to predict future user-item interactions based on historical interactions. While many SR approaches concentrate on user IDs and item IDs, the human perception of the world through multi-modal signals, like text and images, has inspired researchers to delve into constructing SR from multi-modal information without using IDs.
@@ Abstract.
However, the complexity of multi-modal learning manifests in diverse feature extractors, fusion methods, and pre-trained models. Consequently, designing a simple and universal Multi-Modal Sequential Recommendation (MMSR) framework remains a formidable challenge. We systematically summarize the existing multi-modal related SR methods and distill the essence into four core components: visual encoder, text encoder, multimodal fusion module, and sequential architecture. Along these dimensions, we dissect the model designs, and answer the following sub-questions: First, we explore how to construct MMSR from scratch, ensuring its performance either on par with or exceeds existing SR methods without complex techniques. Second, we examine if MMSR can benefit from existing multi-modal pre-training paradigms. Third, we assess MMSR’s capability in tackling common challenges like cold start and domain transferring. Our experiment results across four real-world recommendation scenarios demonstrate the great potential ID-agnostic multi-modal sequential recommendation. Our framework can be found at: https://github.com/MMSR23/MMSR.
Multi-modality, Sequential Recommendation, Transfer Learning
copyright: acmcopyrightjournalyear: 2023doi: XXXXXXX.XXXXXXXprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
Figure 1. Sequential recommendation models the historical interactions between users and items and predicts the next item. Up: IDSR. Every item is represented by an ID Embedding. Bottom: MMSR. Every item is represented by its multi-modality information, such as covers and titles.
1. Introduction
@@
Sequential Recommendation (SR) is a technique that leverages users’ past interactions to recommend their next item of interest. Its implementation in online environments, like e-commerce platforms and streaming media websites, has significantly expanded the range of options available to customers and has attracted increasing attention in the research community. One of the primary challenges in developing SR models is to generate high-quality user and item representations that can enhance recommendation performances.
@@
In mainstream SR scenarios, a user can be represented by all the items that this user has acted upon, with each item encoded as a unique identifier (ID for short), as is shown in the upper part of Figure 1. To provide recommendation, the prevailing paradigm in current community is to maintain a learnable embedding matrix for all item IDs, and item representation in sequences can map as a latent vector by looking up its corresponding ID, we summarize this paradigm as ID-based Sequential Recommendation (IDSR). For example, GRU4Rec (Hidasi et al., 2015) adopts Recurrent Neural Network (RNN) as the sequence encoder, NextItNet (Yuan et al., 2019) adopts Convolutional Neural Network (CNN) as the sequence encoder, SASRec (Kang and McAuley, 2018) adopts unidirectional Transformer (Vaswani et al., 2017) as the sequence encoder, BERT4Rec (Sun et al., 2019) adopts bidirectional Transformer (Devlin et al., 2018) as the sequence encoder. Although straightforward and easy to implement, IDSR has several significant limitations. First, ID representations learned from one domain cannot be transferred across platforms, since the item vocabularies from two non-overlapping domains cannot be shared (Kang et al., 2020; Yuan et al., 2023). Second, IDSR can only learn the representations from the item IDs, leading to recommend cold items from highly skewed data distribution intractable (Wei et al., 2021; Zhu et al., 2021).
Alternatively, we can represent each item not by its ID, but by the multi-modality information (e.g., covers and titles) of the item, which we refer to as Multi-Modality based Sequential Recommendation (MMSR). Recently, several studies begin to utilize intrinsic modality or representations to glean item characteristics and learn user behaviors. For example, ZeroRec (Ding et al., 2021), UniSRec (Hou et al., 2022b), , VQRec (Hou et al., 2022a), MoRec (Yuan et al., 2023) and (Li et al., 2023a; Fu et al., 2023) pioneered the substitution of text or image representations for ID representations, achieving transferable sequential recommendations. TransRec (Wang et al., 2022) and MSM4SR (Zhang et al., 2023b) introduced multi-modal information to yield universal item and user representations. Compared with IDSR, MMSR is naturally endowed with strong transferability, since MMSR can process different items from different domains and platforms via the multi-modal contents of the items without the need for shareable item IDs. Furthermore, MMSR can alleviate the cold-start issue, since text and vision contents of the items provide richer semantics that are otherwise unobserved in the explicitly ID-based interactions.
@@
We recap the basic idea of MMSR methods, which is to extract the multi-modal item embeddings from the pre-trained multi-modal item encoders ( an item encoder usually consists of a visual encoder, a textual encoder, and a multi-modal fusion module) and then adopt a sequence encoder (SASRec (Kang and McAuley, 2018)) to process the multi-modal item embeddings for recommendation training. On top of this backbone, these works design additional techniques to achieve better recommendation performance, such as enhancing the quality of text representations via parametric whitening and mixture-of-experts enhanced adaptor (Hou et al., 2022a), and achieving embedding alignment via contrastive learning with hard negatives (Hou et al., 2022b). Though effective, these techniques are not generalized enough to fit different types of multi-modal item encoders and sequence encoders and usually require careful tuning and selecting of hyperparameters. Meanwhile, the technical breakthroughs in NLP, CV, and RecSys are always evolving, with the new state-of-the-art text encoders (Devlin et al., 2018; Liu et al., 2019; Clark et al., 2020), vision encoders (He et al., 2016; Huang et al., 2020; He et al., 2022), and sequence encoders (Hidasi et al., 2015; Yuan et al., 2019; Kang and McAuley, 2018; Sun et al., 2019) emerging frequently. Excessive pursuit of the additional techniques leads to poor compatibility with different types of multi-modal item encoders and sequence encoders, making it difficult to utilize the technical breakthroughs in the NLP, CV, and RecSys communities.
@@
To address this issue, this paper builds a simple and universal MMSR framework, which can support various types of multi-modal item encoders and sequence encoders and shows performance better or on par with the state-of-the-art multi-modal sequential recommenders. The core idea is to design a plug-and-play MMSR backbone consisting of multi-modal item encoders, multi-modal fusion modules, and the sequence encoder, such that different types of components can be seamlessly integrated into or replaced from the backbone. In our empirical study of MMSR, we try to answer the following key questions: Q1: How can we intuitively build MMSR from scratch, ensuring its performance on par with or even exceeds contemporary SR methods, all without the crutch of complex tricks? Q2: How can we build an MMSR framework such that it can directly benefit from the advances in vision-and-language communities? Q3: Will MMSR be supportive in addressing critical issues in the RecSys community, e.g., cold-start problems and cross-domain transferring?
@@
To answer Q1, we conduct experiments on four large-scale industrial datasets under three SR architectures (i.e., SASRec (Kang and McAuley, 2018), GRURec (Hidasi et al., 2015), and NextItNet (Yuan et al., 2019)), and extensively test the performance of MMSR by equipping it with three TE, three VE and ten fusion strategies (including 4 vanilla coarse-grained, 4 enhanced coarse-grained and 2 fine-grained approaches). To answer Q2, we explore whether MMSR benefits from advances in existing Vision-Language Pre-trained (VLP) models by experimenting with different VLP architectures (e.g., CLIP (Radford et al., 2021) & GroupViT (Xu et al., 2022), ViLT (Kim et al., 2021) & VisualBERT (Li et al., 2019), and FLAVA (Singh et al., 2022)) as well as various representations of vision in VLP (i.e., region-based (Faster, 2015), grid-based (He et al., 2016) and patch-based (Radford et al., 2021)). To answer Q3, we put IDSR and MMSR in the cold-item scenario to evaluate their item representation ability and conduct transferring experiments on two source domains and two target domains, with the source and target datasets being scaled. In addition to exploring the above issues, practical guidance on how to effectively train MMSR would also be provided, e.g., whether full-finetuning the item encoders is necessary, what the impact of information volume brings (i.e., title length of sentences and mask ratio of images), and hyperparameter of different modules (see Appendix F). Our contributions can be outlined as follows:
@@
2. Related Works
2.1. ID-based Sequential Recommendation
Currently, the dominant paradigm of recommender systems is still IDSR, which assigns unique item/user IDs with one-hot embedding. There are thousands of recommenders based solely on IDs as input (Linden et al., 2003; Koren et al., 2009; Rendle, 2010; He et al., 2017; Hidasi et al., 2015). In this paper, we mainly focus on the most advanced and robust sequential recommenders, including RNN-based GRU4Rec (Hidasi et al., 2015), CNN-based NextItNet (Yuan et al., 2019), unidirectional Transformer-based SASRec (Kang and McAuley, 2018) and BERT4Rec (Sun et al., 2019). Although straightforward to implement, IDSR shows limited transferability and inferior performances under the cold-start setting.
2.2. Multi-Modal Sequential Recommendation
Some research focuses on improving recommendations by merging multi-modal content in IDSR. FDSA (Zhang et al., 2019) blends text with ID-embeddings using self-attention, and CARCA (Rashed et al., 2022) adopts cross-attention for item features. Yet, their reliance on the ID model limits cross-domain and platform flexibility.
To address this, ID-agnostic multi-modal recommenders have been introduced. UniSRec (Hou et al., 2022b) uses whitening (Su et al., 2021) to bridge the gap between item text and embeddings. VQRec (Hou et al., 2022a) matches item text with discrete index vectors to derive representations. MSM4SR (Zhang et al., 2023b) pioneers multi-modal information with contrastive learning for transfer. TransRec (Wang et al., 2022) leverages hybrid feedback for broader application. MoRec (Yuan et al., 2023) shows that text or vision can compete with ID representations. (Fu et al., 2023) further optimized MoRec, enhancing efficiency with Adapters (Rebuffi et al., 2017), reducing computational costs, and proving comparable performance to end-to-end methods. PMMRec (Li et al., 2023b) suggested an innovative cross-modal contrastive learning objective that enhances next-item prediction.
@@ 2.2. Multi-Modal Sequential Recommendation
Our MMSR framework integrates diverse multi-modal item and sequence encoders, aiming to create a flexible, universally compatible, and efficient system. It outperforms traditional ID-based systems in addressing recommender system challenges like cold starts and transfer learning. A work closely related to ours is MMRec (Zhou et al., 2023), which outlines a clear process for multi-modal recommendation systems. However, it primarily targets GNN-based and general recommender systems, whereas our focus is on multi-modal sequential recommendation systems.
3. Framework
@@
In this section, we introduce several essential modules in the MMSR framework and illustrate them in Figure 2. For an item with textual and visual information in SR, we fine-tune the corresponding pre-trained encoders to learn its text and vision features, followed by a fusion module to combine the two unimodal features into a cross-modal representation. Note, the combination of both text and vision encoders and fusion modules can be replaced by the existing multi-modal Vison-and-Language Pre-trained (VLP) models. Finally, we feed the item representation into a sequential recommendation network to complete the downstream recommendation task.
@@
Figure 2. Illustration of MMSR. We systematically investigate MMSR along multiple dimensions, including vision and text encoders, multi-modality fusion, multi-modal pre-training paradigms, and downstream sequential architectures.
3.1. Text Encoder (TE)
@@
In this paper, we take BERT (Devlin et al., 2018), RoBERTa (Liu et al., 2019) and ELECTRA (Clark et al., 2020) as the TE. There are two ways of using text embedding, i.e., either providing it to multiple text-specific transformers before the fusion or directly feeding it to a multi-modal fusion module. We adopt the former as our default setting, the latter will be discussed while introducing VLPs in Section 6.1.
3.2. Vision Encoder (VE)
@@
Existing visual representations can be roughly divided into three categories: (1) Region Feature. Many traditional VLP models dominantly utilize region features, which are also known as bottom-up features. They are obtained from an off-the-shelf object detector such as Faster R-CNN (Faster, 2015). (2) Grid Feature. Generated by some CNN-based methods, e.g., ResNet50 (He et al., 2016), it avoids very slow region selection operations. (3) Patch or Patch-ViT. It is a commonly used approach that was first introduced by ViT (Dosovitskiy et al., 2020), showing outstanding performance in image classification tasks. In this paper, we first explore ResNet50 (He et al., 2016), MAE (He et al., 2022), and Swin (Liu et al., 2021) as VE in MMSR without VLPs, and then we give a detailed investigation of the three representations in VLPs in Section 6.2.
3.3. Multi-modal Fusion

### v5

@@
Abstract
@@
Sequential Recommendation (SR) aims to predict future user-item interactions based on historical interactions. While many SR approaches concentrate on user IDs and item IDs, the human perception of the world through multi-modal signals, like text and images, has inspired researchers to delve into constructing SR from multi-modal information without using IDs. However, the complexity of multi-modal learning manifests in diverse feature extractors, fusion methods, and pre-trained models. Consequently, designing a simple and universal Multi-Modal Sequential Recommendation (MMSR) framework remains a formidable challenge. We systematically summarize the existing multi-modal related SR methods and distill the essence into four core components: visual encoder, text encoder, multimodal fusion module, and sequential architecture. Along these dimensions, we dissect the model designs, and answer the following sub-questions: First, we explore how to construct MMSR from scratch, ensuring its performance either on par with or exceeds existing SR methods without complex techniques. Second, we examine if MMSR can benefit from existing multi-modal pre-training paradigms. Third, we assess MMSR’s capability in tackling common challenges like cold start and domain transferring. Our experiment results across four real-world recommendation scenarios demonstrate the great potential ID-agnostic multi-modal sequential recommendation. Our framework can be found at: https://github.com/MMSR23/MMSR.
Index Terms:
Multi-modality, Sequential Recommendation, Transfer Learning.
1 Introduction
@@
Sequential Recommendation (SR) leverages users’ past interactions to recommend their next item of interest. Its implementation in online environments, like e-commerce platforms and streaming media websites, has significantly expanded the range of options available to customers and has attracted increasing attention in the research community. One of the primary challenges in developing SR models is to generate high-quality user and item representations that can enhance recommendation performances.
@@
In mainstream SR scenarios, a user can be represented by all the items that this user has acted upon, with each item encoded as a unique identifier (ID for short), as is shown in the upper part of Figure 1. To provide recommendation, the prevailing paradigm in current community is to maintain a learnable embedding matrix for all item IDs, and item representation in sequences can map as a latent vector by looking up its corresponding ID, we summarize this paradigm as ID-based Sequential Recommendation (IDSR). For example, GRU4Rec [1] adopts Recurrent Neural Network (RNN) as the sequence encoder, NextItNet [2] adopts Convolutional Neural Network (CNN) as the sequence encoder, SASRec [3] adopts unidirectional Transformer [4] as the sequence encoder, BERT4Rec [5] adopts bidirectional Transformer [6] as the sequence encoder. Although straightforward and easy to implement, IDSR has several significant limitations. First, ID representations learned from one domain cannot be transferred across platforms, since the item vocabularies from two non-overlapping domains cannot be shared [7, 8]. Second, IDSR can only learn the representations from the item IDs, leading to recommend cold items from highly skewed data distribution intractable [9, 10]. Additionally, IDSR’s inability to leverage advancements in other fields such as Natural Language Processing (NLP) and Computer Vision (CV) restricts the development of large-scale general model applications within recommender systems, unlike in NLP and CV domains.
Figure 1: Sequential recommendation models the historical interactions between users and items and predicts the next item. Up: IDSR. Every item is represented by an ID Embedding. Bottom: MMSR. Every item is represented by its multi-modality information, such as covers and titles.
To address the cold start problem and enhance the quality of recommendations, numerous Sequential Recommendation (SR) studies [11, 12, 13, 14, 15] have integrated (multi-)modality information, such as text and images, as auxiliary features. These features aim to construct more comprehensive item representations by combining them with ID-based representations. Despite this integration, these models remain dependent on IDs, limiting their applicability across different domains. This raises a critical question: Are user and item IDs indispensable for recommender systems? We argue that the modality or multi-modality information of an item, such as covers and titles, can serve as unique and distinguishable representations, as illustrated in Figure 1 (below). As an alternative approach, we propose representing items not by their IDs but by their multi-modality information, a concept we define as Multi-Modality based Sequential Recommendation (MMSR).
Recent advancements in the field have seen a shift towards directly utilizing modality representations to extract item characteristics and understand user behaviors more effectively. Innovations such as ZeroRec [16], UniSRec [17], and VQRec [18] have been at the forefront of replacing ID representations with text representations, facilitating transferable sequential recommendations. Additionally, TransRec [19] and MSM4SR [20] have incorporated multi-modal information to achieve universal representations for both items and users. MMSR distinguishes itself from ID-based Sequential Recommendation (IDSR) through its inherent transferability. By utilizing the multi-modal contents of items, MMSR can seamlessly process items from various domains and platforms without relying on shareable item IDs. Moreover, MMSR addresses the cold start issue effectively by leveraging the text and visual content of items, which provides a richer semantic layer that remains untapped in strictly ID-based interactions.
@@
We recap the basic idea of MMSR methods, which is to extract the multi-modal item embeddings from the pre-trained multi-modal item encoders (an item encoder usually consists of a visual encoder, a textual encoder, and a multi-modal fusion module) and then adopt a sequence encoder (e.g., SASRec [3]) to process the multi-modal item embeddings for recommendation training. On top of this backbone, these works design additional techniques to achieve better recommendation performance, such as enhancing the quality of text representations via parametric whitening and mixture-of-experts enhanced adaptor [18], and achieving embedding alignment via contrastive learning with hard negatives [17]. Though effective, these techniques are not generalized enough to fit different types of multi-modal item encoders and sequence encoders and usually require careful tuning and selecting of hyperparameters. Meanwhile, the technical breakthroughs in NLP, CV, and RecSys are always evolving, with the new state-of-the-art text encoders [6, 21, 22], vision encoders [23, 24, 25], and sequence encoders [1, 2, 3, 5] emerging frequently. Excessive pursuit of the additional techniques leads to poor compatibility with different types of multi-modal item encoders and sequence encoders, making it difficult to utilize the technical breakthroughs in the NLP, CV, and RecSys communities.
@@
To address this issue, this paper builds a simple and universal MMSR framework, which can support various types of multi-modal item encoders and sequence encoders and shows performance better or on par with the state-of-the-art multi-modal sequential recommenders. The core idea is to design a plug-and-play MMSR backbone consisting of multi-modal item encoders, multi-modal fusion modules, and the sequence encoder, such that different types of components can be seamlessly integrated into or replaced from the backbone. In our empirical study of MMSR, we try to answer the following key questions: Q1: How can we intuitively build MMSR from scratch, ensuring its performance on par with or even exceeds contemporary SR methods, all without the crutch of complex tricks? Q2: How can we build an MMSR framework such that it can directly benefit from the advances in vision-and-language communities? Q3: Will MMSR be supportive in addressing critical issues in the RecSys community, e.g., cold-start problems and cross-domain transferring? To answer these questions, we conduct comprehensive empirical studies along multiple dimensions, including text encoders (TE), vision encoders (VE), fusion strategies, multi-modal encoders paradigms, and RS architectures.
@@
To answer Q1, we conduct experiments on four large-scale industrial datasets under three SR architectures (i.e., SASRec [3], GRURec [1], and NextItNet [2]), and extensively test the performance of MMSR by equipping it with three TE, three VE and ten fusion strategies (including 4 vanilla coarse-grained, 4 enhanced coarse-grained and 2 fine-grained approaches). To answer Q2, we explore whether MMSR benefits from advances in existing Vision-Language Pre-trained (VLP) models by experimenting with different VLP architectures (e.g., CLIP [26] & GroupViT [27], ViLT [28] & VisualBERT [29], and FLAVA [30]) as well as various representations of vision in VLP (i.e., region-based [31], grid-based [23] and patch-based [26]). To answer Q3, we put IDSR and MMSR in the cold-item scenario to evaluate their item representation ability and conduct transferring experiments on two source domains and two target domains, with the source and target datasets being scaled. In addition to exploring the above issues, practical guidance on how to effectively train MMSR would also be provided, e.g., whether full-finetuning the item encoders is necessary, what the impact of information volume brings (i.e., title length of sentences and mask ratio of images), and hyperparameter of different modules.
Our contributions can be outlined as follows:
@@
2 Related Works
2.1 ID-based Sequential Recommendation
In the domain of recommender systems, ID-based Sequential Recommendation (IDSR) stands as the dominant paradigm. This approach relies on encoding unique identifiers for items and users via one-hot embeddings, serving as the foundation for numerous recommendation models. Seminal contributions in this area include works such as [32, 33, 34, 35, 1], which have substantially influenced the development of IDSR methodologies. Our discussion further extends to state-of-the-art sequential recommenders, encompassing RNN-based models like GRU4Rec [1], CNN-based models such as NextItNet [2], and Transformer-based frameworks, notably SASRec [3] and BERT4Rec [5].
It is important to highlight that the majority of these SR approaches, albeit innovative, primarily rely on ID features—ranging from item IDs to attribute IDs. This reliance inherently subjects them to the limitations of the cold-start problem and restricts their transferability across different contexts and domains.
2.2 IDSR with Multi-modal Side Features
To address this limitation, several researchers have proposed multi-modality recommenders that integrate auxiliary features to enrich item and user representations.
Incorporating multi-modal content into IDSR, FDSA [11] combined text with ID-embeddings using self-attention, while CARCA [15] employed cross-attention to enhance item features. These methods are limited because they depend on ID-based models, reducing their effectiveness across various domains and platforms. Despite potential benefits, multi-modal recommenders still focus on items, limiting their applicability elsewhere.
2.3 Multi-Modal Sequential Recommendation
Addressing the limitations above, which are constrained by their reliance on unique identifiers for items, researchers introduce a paradigm shift with Multi-Modal Sequential Recommendation (MMSR). Unlike IDSR of side features, which primarily depends on ID information treating modal features as auxiliary, MMSR emphasizes scenarios where only multi-modal information (e.g., texts and images) serves as the input for recommendations, thereby facilitating the development of transferable RecSys by leveraging knowledge transfer across domains and platforms to mitigate data sparsity issues.
The trend towards purely modality-based recommendations is rising, with literature covering scenarios of some unimodal Modality-based SR. For example, ZESRec [16] employs pre-extracted text embeddings as transferable item representations for cross-task knowledge transfer, while UniSRec [17] utilizes whitening techniques [36] to bridge the gap between item text and embeddings, deriving item representations through code embedding tables. VQRec [18] matches item text with discrete index vectors for representation. MoRec [8] showing that purely modality-based models can perform comparably to ID-based models by substituting ID embeddings with text or visual representations, demonstrating improvements in feature extractors for text and image-based SR models in various scenarios. [37] further optimized MoRec, enhancing efficiency with Adapters [38], reducing computational costs, and proving comparable performance to end-to-end methods.
Multi-modal recommender systems have been developing. TransRec [19] optimizes hybrid modal sequences for better item representation. MISSRec [39] incorporates user-adaptive fusion for dynamic attention to modalities. MSM4SR [20] merges item images and textual keywords with visual features via BERT for user behavior analysis. [40] leverages texts, images, and prices with a re-learning strategy for task-specific challenges. [41] enhances feature interactions using an ID-aware Multi-modal Transformer and online distillation. PMMRec [42] employs cross-modal contrastive learning for integrating user behavior into item representations, supporting diverse learning modes.
Our MMSR framework aligns with recent advancements yet addresses the challenge of integrating diverse multi-modal item and sequence encoders. We aim to create a flexible, universally compatible, and efficient MMSR framework that works seamlessly with NLP, CV, and RecSys technologies. It has shown to surpass traditional ID-based systems in tests, proving its effectiveness in solving recommender system issues like cold starts and transfer learning.
@@ 2.3 Multi-Modal Sequential Recommendation
A work closely related to ours is MMRec [43], which outlines a clear process for multi-modal recommendation systems. However, it primarily targets GNN-based and general recommender systems, whereas our focus is on multi-modal sequential recommendation systems.
3 Framework
@@
In this section, we introduce essential modules in the MMSR framework and illustrate them in Figure 2. For an item with textual and visual information in SR, we fine-tune the corresponding pre-trained encoders to learn its text and vision features, followed by a fusion module to combine the two unimodal features into a cross-modal representation. Note, the combination of both text and vision encoders and fusion modules can be replaced by the existing multi-modal Vison-and-Language Pre-trained (VLP) models. Finally, we feed the item representation into a sequential recommendation network to complete the downstream recommendation task.
@@
Figure 2: Illustration of MMSR. We systematically investigate MMSR along multiple dimensions, including vision and text encoders, multi-modality fusion, multi-modal pre-training paradigms, and downstream sequential architectures.
3.1 Text Encoder (TE)
@@
In this paper, we take BERT [6], RoBERTa [21] and ELECTRA [22] as the TE. There are two ways of using text embedding, i.e., either providing it to multiple text-specific transformers before the fusion or directly feeding it to a multi-modal fusion module. We adopt the former as our default setting, the latter will be discussed while introducing VLPs in Section 6.1.
3.2 Vision Encoder (VE)
@@
Existing visual representations can be roughly divided into three categories: (1) Region Feature. Many traditional VLP models dominantly utilize region features, which are also known as bottom-up features. They are obtained from an off-the-shelf object detector such as Faster R-CNN [31]. (2) Grid Feature. Generated by some CNN-based methods, e.g., ResNet50 [23], it avoids very slow region selection operations. (3) Patch or Patch-ViT. It is a commonly used approach that was first introduced by ViT [44], showing outstanding performance in image classification tasks. In this paper, we first explore ResNet50 [23], MAE [25], and Swin [45] as VE in MMSR without VLPs, and then we give a detailed investigation of the three representations in VLPs in Section 6.2.
3.3 Multi-modal Fusion